SOLVING THE SEXTIC BY ITERATION: A STUDY IN COMPLEX GEOMETRY AND

DYNAMICS

SCOTT CRASS

Abstract. We use the Valentiner action of A_{6} on ℂ ℙ^{2} to develop an iterative algorithm for the solution of the general sextic equation over ℂ, analogous to Doyle and McMullen’s algorithm for the quintic.